009 (part 1 of 3) 10.0 points A 4.8 kg object is subjected to two forces, F(2.3 N) i (-2.6 N) j and F (4.9 N) +(-11.9 N) . The object is at rest at the origin at time t -0. What is the magnitude of the objects ac- celeration? Answer in units of m/s?. 010 (part 2 of 3) 10.0 points What is the magnitude of the velocity at t- 3.8 s ? Answer in units of m/s. 011 (part 3 of 3) 10.0 points What is the magnitude of the objects position at t 3.8 s ? Answer in units of m 7:25 P
Answer #1

Given F, (2.3i _ 2.67)N

E, (4.92-11.97)N

--------------

Net Force =mass *acceleration

F1 + F2-mã

(2.3i-2.6j)N+(4.9i-11.9j)N=4.8kg*vec{a}

(7.22-14.5) = 4.8 *a

(7.2i-14.5j)/4.8=vec{a}

m s

a - /1.52 (-3.021)2 CU

ANSWER: a 3.373ms (l

=======================

Part 2

initially object is at rest

Use Formula v=u+at

v=0m/s+3.373m/s^{2}*3.8s

v=0+3.373*3.8

ANSWER: v- 12.8174m/s

===========================

Part 3

Use Formula s=ut+1/2at^{2}

initially object is at rest

s=1/2at^{2}

s=0.5*3.373m/s^{2}*(3.8s)^{2}

-053373 *3,82

ANSWER: {color{Red} s=24.35m}

=====================